ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 183, 2011

Abdenur, F., C. Bonatti and S. Crovisier Nonuniform hyperbolicity for C^{1}-generic diffeomorphisms

1

Ben Yaacov, I., A. Berenstein and C. W. Henson Modeltheoretic independence in the Banach lattices L_{p}(μ)

285

Berenstein, A. See I. Ben Yaacov Blau, H. I. See B. Xu Bonatti, C. See F. Abdenur

Bowen, L. Weak isomorphisms between Bernoulli shifts  93

Carnovale, G. and J. Cuadra On the subgroup structure of the full Brauer group of Sweedler Hopf algebra

61

Cimprič, J., M. Marshall and T. Netzer Closures of quadratic modules

445

Crovisier, S. See F. Abdenur Cuadra, J. See G. Carnovale Danilenko, A. I. and A. del Junco Almost continuous orbit equiv-

alence for non-singular homeomorphisms

165

Domingo-Juan, M. C., A. Lluch and V. Miquel Upper bounds for the first Dirichlet eigenvalue of a tube around an algebraic complex curve of ℂP^{1}(λ)

189

Figiel, T., W. B. Johnson and A. Pełczyński Some approximation properties of Banach spaces and Banach lattices

199

García, G. A. and A. García Iglesias Finite-dimensional pointed Hopf algebras over 𝕊_{4}

417

García Iglesias, A. See G. A. García

Greschonig, G. Nilpotent extensions of Furstenberg transformations  381
Hazrat, R. and A. R. Wadsworth SK_{1} of graded division algebras  117

Henson, C. W. See I. Ben Yaacov

Holst, T. and B. Shapiro On higher Heine–Stieltjes polynomials  321

Huang, W., P. Lu and X. Ye Measure-theoretical sensitivity and equicontinuity

233

Jiménez, C. H. and M. Naszódi On the extremal distance between two convex bodies

103

Johnson, W. B. See T. Figiel del Junco, A. See A. I. Danilenko Lluch, A. See M. C. Domingo-Juan Lu, P. See W. Huang Marshall, M. See J. Cimprič Miquel, V. See M. C. Domingo-Juan Naszódi, M. See C. H. Jiménez Netzer, T. See J. Cimprič Pełczyński, A. See T. Figiel Pusti, S. and R. P. Sarkar Spectral analysis for radial sections of

some homogeneous vector bundles on certain noncompact Riemannian symmetric spaces

475

Sarkar, R. P. See S. Pusti Shapiro, B. See T. Holst Shen, Y. On Grunsky inequalities and the exact domain of variability

of the Grunsky functional

399

Wadsworth, A. R. See R. Hazrat Xu, B. and H. I. Blau On pseudocyclic table algebras and applica-

tions to pseudocyclic association schemes

347

Ye, X. See W. Huang